# Quasi-likelihood inference for SDE with mixed-effects observed at high frequency

### v1

@@ 3.1. Diffusion parameter: profile quasi-likelihood
In this first step, we regard the drift components in (2.1) as nuisance elements in estimating the target parameter $\vartheta_{1}$. Recall we are using the notation:
@@
Recall the notation $S_{i,j-1}(\eta)=c_{i,j-1}^{2}(\eta)$. If we temporarily regard the random-effect parameters $\{\tau_{i}\}_{i}$ as statistical parameters, then it is natural from (3.2) to consider the following Gaussian quasi-likelihood:
$(\eta,\tau_{1},\dots,\tau_{N})\mapsto\sum_{i,j}\log\phi\big{(}Y_{ij};Y_{i,j-1},h\tau_{i}\,S_{i,j-1}(\eta)\big{)}.$
@@
$\widehat{\tau}_{i}(\eta):=\frac{1}{n}\sum_{j}S^{-1}_{i,j-1}(\eta)\,y_{ij}^{2},\qquad i=1,\dots,N.$ (3.3)
If $\{\tau_{i}\}_{i}$ were observed in addition to $\{Y_{ij}\}_{i,j}$, then taking the $\tau_{i}$-conditional approximation (3.2) into account we may define the quasi-log-likelihood as follows:
@@
Ignoring the irrelevant terms free from $\vartheta_{1}$ in (3.5), we define the following random function which we call the first-stage profile quasi-likelihood:
@@
$\mathbb{H}_{11,N,n}(\eta)$ $:=-\frac{1}{2}\sum_{i}\bigg{\{}\sum_{j}\log S_{i,j-1}(\eta)+n\log\bigg{(}\frac{1}{n}\sum_{j}S_{i,j-1}^{-1}(\eta)\,y_{ij}^{2}\bigg{)}\bigg{\}},$ (3.7)
@@
  1. (1)
    The random function $\mathbb{H}_{11}(\eta)$ based on (3.3) is similar to the GQLF for a discretely observed diffusion process with an unknown time scale [13], where we considered profiling out the time scale parameter from the conventional Gaussian quasi-likelihood and derived the asymptotic properties of the associated estimators. Indeed, our technical arguments in this paper partly owe those in [13], while our situation is more complicated since we have many processes. We also note that the linear dependence of $\tau_{i}$ in the drift coefficient of (2.1) is never essential: indeed, it does not matter how $\tau_{i}$ depends on the drift coefficient in the second-stage drift estimation where we will plug in $\widehat{\tau}_{i}$ as a known quantity.
  2. (2)
@@
    to estimate $\theta_{\tau}$ based on the statistics $\widehat{\tau}^{\ast}_{i}:=n^{-1}\sum_{j}S^{-1}_{i,j-1}\,y_{ij}^{2}$.
@@
$Y_{i}(t)$ $=Y_{i}(0)+\int_{0}^{t}\mu_{i}(s)ds+\sqrt{\tau_{i}}\int_{0}^{t}c_{i}(s;\eta)dw_{i}(s),$
where $\mu_{i}$ are unknown $(\mathcal{F}_{i,t})$-adapted processes. This corresponds to a sort of population modeling of the non-ergodic volatility regression [26]. This section’s contents remain valid under mild conditions on $\mu_{i}$.
@@ 3.2. Drift parameter: marginal quasi-likelihood
Having obtained the first-stage estimate $\widehat{\vartheta}_{1}=(\widehat{\eta},\widehat{\theta}_{\tau})$, we proceed to estimating the remaining drift parameter $\vartheta_{2}=(\mu,\Sigma_{r})$. The Euler approximation in this second stage takes the drift term into account:
@@
Recall that we are assuming that $\varphi_{i}\sim\text{i.i.d.}\leavevmode\nobreak\ N_{p_{\varphi}}(\mu,\Sigma)$, which is truly degenerate: recall (2.4) and (2.5).
@@
As in [6] and [5], we temporarily integrate out the variable $\varphi_{1},\dots,\varphi_{N}$ with respect to the Gaussian distribution $N_{p_{\varphi}}(\mu,\Sigma)^{\otimes N}$ as if $\Sigma=\mathrm{diag}(O,\Sigma_{r})$ is non-degenerate: for each $i\leq N$, some matrix manipulations and the integration of the Gaussian density give
@@
are statistics. The right-hand side of (3.15) is valid for the degenerate $\Sigma$ as soon as $\widehat{M}_{i}$ is invertible, which is true with probability tending to $1$ for $n\to\infty$ under Assumption 2.2; more precisely, one may go through the above calculations to obtain (3.15) first by replacing $\Sigma$ by $\Sigma+m^{-1}I_{p_{\varphi}}$ for $m>0$ and then applying Scheffé’s lemma for $m\to\infty$.
Taking the product of (3.15) for $i=1,\dots,N$ and then taking the logarithm, we arrive at the following explicit Gaussian quasi-likelihood for estimating $\vartheta_{2}$:
@@
$=-\frac{p_{\varphi}N}{2}\log(2\pi)-\frac{1}{2}\sum_{i}\log\det\big{(}\widehat{M}_{i}^{-1}+\Sigma\big{)}$
${}\qquad-\frac{1}{2}\sum_{i}(\widehat{M}_{i}^{-1}\widehat{v}_{i}-\mu)^{\top}(\widehat{M}_{i}^{-1}+\Sigma)^{-1}(\widehat{M}_{i}^{-1}\widehat{v}_{i}-\mu).$
(Here and in what follows, for brevity, we often keep using the notation $\Sigma=\mathrm{diag}(O,\Sigma_{r})$) Thus, we obtained the simple Gaussian form (3.17), ignoring several terms in computing (3.15). We can define an estimator of $\vartheta_{2}$ (of $(\mu,{\rm vech}(\Sigma_{r}))$) by any element
$\widehat{\vartheta}_{2}=(\widehat{\mu},\widehat{\Sigma}_{r})\in\mathop{\rm argmax}\mathbb{H}_{2}.$ (3.18)
We note that $\mathbb{H}_{2}(\vartheta_{2})$ is formally the same as in the no-truncated version of the random function $\mathbf{W}_{n}(X_{i,n},\bm{\mu},\bm{\Omega})$ in [5, Eq.(31)].
@@
$M_{i,T}:=\tau_{i}\int_{0}^{T}\frac{a_{i}^{\otimes 2}(t)}{S_{i}(t)}dt,\qquad\widehat{M}_{i}:=\widehat{\tau}_{i}\,h\sum_{j}\widehat{S}_{i,j-1}^{-1}a_{i,j-1}^{\otimes 2}.$ (3.19)
@@
$\mathbb{Y}_{2,0}(\vartheta_{2})$ $:=E\left[\log\left(\frac{\phi_{p_{\varphi}}\left(M_{1}^{-1}v_{1};\,\mu,\,M_{1}^{-1}+\Sigma\right)}{\phi_{p_{\varphi}}\left(M_{1}^{-1}v_{1};\,\mu_{0},\,M_{1}^{-1}+\Sigma_{0}\right)}\right)\right].$ (3.20)
The next technical assumption will be used to estimate the drift parameters.
@@
  1. (1)
    There exist positive constants $\varepsilon_{1},\varepsilon_{2}$ for which
    $\min_{i\leq N}\left(\lambda_{\min}(M_{i})^{-2-\varepsilon_{1}}+\lambda_{\min}(\widehat{M}_{i})^{-1-\varepsilon_{2}}\right)=O_{p}(1).$ (3.21)
  2. (2)
@@
Simplification of 3.5 would be possible in a similar manner to the proof of [13, Lemma 5.2] to handle the inverse of $\widehat{\tau}_{i}$; still, we need to be careful since we are now having the random effect parameters $(\tau_{i},\varphi_{i})$.
As in [5], we introduce the following notation:
$B_{i,T}(\Sigma_{r})$ $=(M_{i,T}^{-1}+\Sigma)^{-1},$
$A_{i,T}(\mu,\Sigma_{r})$ $=B_{i,T}(\Sigma_{r})\big{(}M_{i,T}^{-1}U_{i,T}-\mu\big{)},$
@@
Finally, let $\mathcal{J}_{2,0}\in\mathbb{R}^{q_{\varphi}}$ ($q_{\varphi}:=p_{\varphi}+p_{\varphi}(p_{\varphi}+1)/2$) denote the covariance matrix associated with the random variable
$\begin{pmatrix}\tau_{1}^{-1}A_{1,T}(\mu,\Sigma_{r})\\ \frac{1}{2}\tau_{1}^{-1}A_{1,T}(\mu,\Sigma_{r})^{\otimes 2}-B_{1,T}(\Sigma_{R})\end{pmatrix}$. (3.23)
@@
Here is a remark on yet another estimation procedure. Numerical optimisation (3.18) can be hard if $p_{\varphi}$ is large. We may attempt to split the estimation stages further in a more easy-to-compute manner. Specifically, given any symmetric positive definite $p_{\varphi}\times p_{\varphi}$-matrix $M$, we define an auxiliary estimator $\widehat{\mu}_{0}(M)$ of $\mu_{0}$ by any element in $\mathop{\rm argmax}\mathbb{H}_{2}(\cdot,M)$. Most simply, we consider $M=I_{p_{\varphi}}$. Then, the first-order condition $\partial_{\mu}\mathbb{H}_{2}(\widehat{\mu}_{0},I_{p_{\varphi}})=0$ gives the explicit estimator:
$\widehat{\mu}_{0}(I_{p_{\varphi}})=\left(\sum_{i}(\widehat{M}_{i}^{-1}+I_{p_{\varphi}})^{-1}\right)^{-1}\sum_{i}(\widehat{M}_{i}^{-1}+I_{p_{\varphi}})^{-1}\widehat{M}_{i}^{-1}\widehat{v}_{i}.$ (3.24)
We then proceed with the estimation of $\Sigma$ by any element
$\widehat{\Sigma}_{0}\in\mathop{\rm argmax}\mathbb{H}_{2}(\widehat{\mu}_{0},\cdot),$ (3.25)
which requires numerical optimization. Write $\widehat{\vartheta}_{2,0}:=(\widehat{\mu}_{0},\widehat{\Sigma}_{0})$. To gain the asymptotic efficiency, we can resort to the one-step estimator of the Newton-Raphson type:
$\widehat{\vartheta}_{2,1}=(\widehat{\mu}_{1},\widehat{\Sigma}_{1})$ $:=\widehat{\vartheta}_{2,0}-\left(\partial_{\vartheta_{2}}^{2}\mathbb{H}_{2}\big{(}\widehat{\vartheta}_{2,0}\big{)}\right)^{-1}\,\partial_{\vartheta_{2}}\mathbb{H}_{2}\big{(}\widehat{\vartheta}_{2,0}\big{)}.$ (3.26)
This alternating procedure would be of practical value.
Remark 3.8.
In the current model setup, the random effect parameters in the drift coefficients take the multiplicative form $\tau_{i}\varphi_{i}$, where $\tau_{i}$ and $\varphi_{i}$ are mutually independent. However, because of the stepwise nature of our inference procedure, we could flexibly modify this point in several ways; for example, we could incorporate $\varphi_{i}$ instead of $\tau_{i}\varphi_{i}$ with the conditional probability structure $\mathcal{L}(\varphi_{i}|\tau_{i})=N_{p_{\varphi}}(\mu,\tau_{i}\Sigma)$ as in [6] and [5]. The modifications are relatively straightforward, and we will not delve into the details here.
@@ 3.3. Connections to existing results
It is worth noting that the convergence rates established in our theoretical analysis are in line with several well-known results in the literature even though these results were originally obtained under (quasi-)maximum likelihood estimation procedures—which is not exactly the case in our framework. In classical nonlinear mixed-effects models, it is established that, when $N$ and $n$ both tend to infinity fixed effects converge at rate $\sqrt{Nn}$, while parameters governing the distribution of random effects converge at rate $\sqrt{N}$, where $N$ denotes the number of individuals and $n$ the number of observations per individual [22]. In more specialized settings, which correspond to particular cases of our model, it has been shown that, for $T$ fixed and $N,n$ tending to infinity, fixed effects in the diffusion coefficient also converge at rate $\sqrt{Nn}$, while fixed effects in the drift and the parameters of the random effects distributions converge at rate $\sqrt{N}$ [7, 8, 5, 6]. These findings are consistent with our results, and further supported by the structure of the asymptotic covariance matrix for the estimators of the diffusion parameters, which reveals the same level of precision comparable to that obtained if the individual-level random effects were observed.
@@ 4. Numerical experiments
$\begin{split}dY_{i}(t)&=-\frac{\tau_{i}\varphi_{i}}{\sqrt{1+Y_{i}(t)^{2}}}dt+\sqrt{\tau_{i}}\exp(\eta t/2)dW_{i}(t)\\ \tau_{i}&\sim\textrm{LogNormal}(\alpha,\sigma),\;\varphi_{i}\sim\mathcal{N}(\mu,\omega^{2}),\;i=1,\ldots,N,\end{split}$
@@
$\begin{split}dY_{i}(t)&=\tau_{i}\left(-\frac{\varphi_{i1}Y_{i}(t)}{\sqrt{1+Y_{i}(t)^{2}}}-\frac{\varphi_{i2}}{\sqrt{1+Y_{i}(t)^{2}}}\right)dt+\sqrt{\tau_{i}}\exp(\eta\textrm{atan}(Y_{i}(t)))dW_{i}(t)\\ \tau_{i}&\sim\textrm{Weibull}(\alpha,\lambda),\;\varphi_{i1}\sim\mathcal{N}(\mu_{1},\omega_{1}^{2}),\;\varphi_{i2}:=\mu_{2},\;i=1,\ldots,N,\end{split}$
@@
$\begin{split}dY_{i}(t)&=\tau_{i}\left(-\frac{\varphi_{i1}Y_{i}(t)}{\sqrt{1+Y_{i}(t)^{2}}}-\frac{\varphi_{i2}}{\sqrt{1+Y_{i}(t)^{2}}}\right)dt+\sqrt{\tau_{i}}\exp(\eta\textrm{atan}(Y_{i}(t)))dW_{i}(t)\\ \tau_{i}&\sim\mathrm{Exp}(\lambda),\;(\varphi_{i1},\varphi_{i2})^{\top}\sim\mathcal{N}((\mu_{1},\mu_{2})^{\top},\Omega),\;i=1,\ldots,N.\end{split}$
For each model, data sets are simulated according to different scenarios that combine several values for $N$ (in $\{200,500\}$), several values for $T$ (in $\{5,10\}$) and several time steps $h$ (in $\{0.005,0.001\}$) between consecutive observations leading to different numbers of observations per trajectory $n$. Each observation path is simulated according to its Euler scheme with a very small time step $\delta=0.0001$. In each scenario, $500$ datasets are simulated and the model parameters are estimated using the procedures described above. To illustrate the theoretical results obtained above, the parameters $\theta_{\tau}$ are also estimated from the simulated $\tau_{i}$’s as if they were directly observed. The true parameter values used in this numerical study are:
  • Model 1: $\eta=0.5$, $\alpha=-0.7$, $\sigma=0.7$, $\mu=2$, $\omega^{2}=1$;
  • Model 2: $\eta=0.5$, $\alpha=1$, $\lambda=0.6$, $\mu_{1}=2$, $\mu_{2}=1$, $\omega_{1}^{2}=1$,
  • Model 3: $\eta=0.5$, $\lambda=1$, $\mu_{1}=2$, $\mu_{2}=1$, $\Omega=\begin{pmatrix}1&-0.2\\ -0.2&0.5\end{pmatrix}$.
The results are depicted in Tables  1-2 (Figure 1) for Model 1, Tables 3-4 (Figure 2) for Model 2 and Tables 5-6 (Figure 3) for Model 3. Each table summarizes the empirical means and standard deviations of the estimates obtained from the $500$ simulated data sets in each of the $(N,n,T)$ configurations.
@@
Across these 3 models, the good results obtained in this simulation study show that the estimation procedure proposed here is sufficiently generic to handle different distributions for random effects in the diffusion coefficient, fixed and random effects in drift, and a more or less complex covariance structure between random drift effects. The ability to handle any distribution with positive support for the random effect in the diffusion coefficient is a real advance over the state of the art.
Table 1. Simulations under Model 1: empirical mean and empirical standard error (in brackets) of $500$ the diffusion parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\eta=0.5$, $\alpha=-0.7$, $\sigma=0.7$. $\widehat{\eta_{y}}$, $\widehat{\alpha_{y}}$ and $\widehat{\sigma_{y}}$ stand for the estimates computed from the trajectories observations, whereas $\widehat{\alpha_{\tau}}$ and $\widehat{\sigma_{\tau}}$ denote the estimates computed from the simulated $\tau_{i}$’s.
$(N,T)$ $n$ $\widehat{\eta_{y}}$ $\widehat{\alpha_{y}}$ $\widehat{\sigma_{y}}$ $\widehat{\alpha_{\tau}}$ $\widehat{\sigma_{\tau}}$
@@
Table 2. Simulations under Model 1: empirical mean and empirical standard error (in brackets) of $500$ the drift parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\mu=2$, $\omega^{2}=1$.
$(N,T)$ $n$ $\widehat{\mu}$ $\widehat{\omega^{2}}$
@@
Table 3. Simulations under Model 2: empirical mean and empirical standard error (in brackets) of $500$ the diffusion parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\eta=0.5$, $\alpha=1$, $\lambda=0.6$. $\widehat{\eta_{y}}$, $\widehat{\lambda_{y}}$ and $\widehat{\alpha_{y}}$ stand for the estimates computed from the trajectories observations, whereas $\widehat{\lambda_{\tau}}$ and $\widehat{\alpha_{\tau}}$ denote the estimates computed from the simulated $\tau_{i}$’s.

### v2

@@ 3.1. Diffusion parameter: profile quasi-likelihood
In this first step, we regard the drift components in (2.1) as nuisance elements in estimating the target parameter $\vartheta_{1}$. Recall that we are using the notation:
@@
Recall the notation $S_{i,j-1}(\eta)=c_{i,j-1}^{2}(\eta)$. If we temporarily regard the random-effect parameters $\{\tau_{i}\}_{i}$ as statistical (non-random) parameters, then it is natural from (3.2) to consider the following Gaussian quasi-likelihood:
$(\eta,\tau_{1},\dots,\tau_{N})\mapsto\sum_{i,j}\log\phi\big(\overline{y}_{ij};Y_{i,j-1},h\tau_{i}\,S_{i,j-1}(\eta)\big).$
@@
$\widehat{\tau}_{i}(\eta):=\frac{1}{n}\sum_{j}S^{-1}_{i,j-1}(\eta)\,\overline{y}_{ij}^{2},\qquad i=1,\dots,N.$ (3.3)
If $\{\tau_{i}\}_{i}$ were observed in addition to $\{\overline{y}_{ij}\}_{i,j}$, then taking the $\tau_{i}$-conditional approximation (3.2) into account we may define the quasi-log-likelihood as follows:
@@
Ignoring the irrelevant terms free from $\vartheta_{1}$ in (3.5), we define the following random function, which we call the first-stage profile quasi-likelihood:
@@
$\mathbb{H}_{11,N,n}(\eta)$ $:=-\frac{1}{2}\sum_{i}\bigg\{\sum_{j}\log S_{i,j-1}(\eta)+n\log\bigg(\frac{1}{n}\sum_{j}S_{i,j-1}^{-1}(\eta)\,\overline{y}_{ij}^{2}\bigg)\bigg\},$ (3.7)
@@
  1. (i)
    The random function $\mathbb{H}_{11}(\eta)$ based on (3.3) is similar to the GQLF for a discretely observed diffusion process with an unknown time scale [EguMas19], where we considered profiling out the time scale parameter from the conventional Gaussian quasi-likelihood and derived the asymptotic properties of the associated estimators. Indeed, our technical arguments in this paper partly owe to those in [EguMas19], while our situation is more complicated because of the presence of the random-effect parameters. We also note that the linear dependence of $\tau_{i}$ in the drift coefficient of (2.1) is never essential: indeed, it does not matter how $\tau_{i}$ depends on the drift coefficient in the second-stage drift estimation, where we will plug in $\widehat{\tau}_{i}$ as a known quantity.
  2. (ii)
@@
    to estimate $\theta_{\tau}$ based on the statistics $\widehat{\tau}^{\ast}_{i}:=n^{-1}\sum_{j}S^{-1}_{i,j-1}\,\overline{y}_{ij}^{2}$.
@@
$Y_{i}(t)$ $=Y_{i}(0)+\int_{0}^{t}\mu_{i}(s)ds+\sqrt{\tau_{i}}\int_{0}^{t}c(Y_{i}(s);\eta)dw_{i}(s),$
where $\mu_{i}$ are unknown $(\mathcal{F}_{i,t})$-adapted processes. This corresponds to a sort of population modeling of the non-ergodic volatility regression [UchYos13]. The claims in Theorem 3.2 remain valid under mild conditions on $\mu_{i}$.
@@ 3.2. Drift parameter: marginal quasi-likelihood
Having obtained the first-stage estimate $\widehat{\vartheta}_{1}=(\widehat{\eta},\widehat{\theta}_{\tau})$, we proceed to estimating the remaining drift parameter $\vartheta_{2}=(\varphi_{f},\Sigma_{r})$. The Euler approximation in this second stage takes the drift term into account:
@@
We are assuming that $\varphi_{i}=(\varphi_{f},\varphi_{r,i})\sim\text{i.i.d.}\penalty 10000\ N_{p_{\varphi}}(\mu,\Sigma)$, which is truly degenerate as soon as we have the fixed-effect parameter $\varphi_{f}$ while $\varphi_{r,i}$ is assumed to be non-degenerate: recall (2.4) and (2.5).
@@
As in [delattre2018parametric] and [DelGenLar18-2], we integrate out the variable $\varphi_{1},\dots,\varphi_{N}$ with respect to the Gaussian distribution $N_{p_{\varphi}}(\mu,\Sigma)^{\otimes N}$ as if $\Sigma=\mathrm{diag}(O,\Sigma_{r})$ is non-degenerate: for each $i\leq N$, some matrix manipulations and the integration of the Gaussian density give
@@
are statistics; note that $\widehat{M}_{i}$ is a.s. non-negative definite. The right-hand side of (3.15) is valid for the degenerate $\Sigma$ as soon as $\widehat{M}_{i}$ is positive definite, which is true with probability tending to $1$ for $n\to\infty$ under Assumption 2.2; specifically, one may go through the above calculations to obtain (3.15) first by replacing $\Sigma$ by $\Sigma+m^{-1}I_{p_{\varphi}}$ for $m>0$ and then applying Scheffé’s lemma for $m\to\infty$.
Taking the product of (3.15) for $i=1,\dots,N$ and then taking the logarithm, we arrive at the following second-stage marginal quasi-likelihood for estimating $\vartheta_{2}$:
@@
$=-\frac{p_{\varphi}N}{2}\log(2\pi)-\frac{1}{2}\sum_{i}\log\det\big(\widehat{M}_{i}^{-1}+\mathrm{diag}(O,\Sigma_{r})\big)$
${}\qquad-\frac{1}{2}\sum_{i}(\widehat{M}_{i}^{-1}+\mathrm{diag}(O,\Sigma_{r}))^{-1}\left[\left(\widehat{M}_{i}^{-1}\widehat{v}_{i}-(\varphi_{f},0)^{\top}\right)^{\otimes 2}\right].$
Then, we define an estimator of $\vartheta_{2}$ by any element
$\widehat{\vartheta}_{2}=(\widehat{\varphi}_{f},\widehat{\Sigma}_{r})\in\mathop{\rm argmax}\mathbb{H}_{2}.$ (3.18)
We note that $\mathbb{H}_{2}(\vartheta_{2})$ is formally the same as in the no-truncated version of the random function $\mathbf{W}_{n}(X_{i,n},\bm{\mu},\bm{\Omega})$ in [DelGenLar18-2, Eq.(31)].
@@
$M_{i}=M_{i,T}:=\tau_{i}\int_{0}^{T}\frac{a_{i}^{\otimes 2}(t)}{S_{i}(t)}dt,\qquad v_{i}=v_{i,T}:=\int_{0}^{T}\frac{a_{i}(t)}{S_{i}(t)}dY_{i}(t).$ (3.19)
@@
$\mathbb{Y}_{2,0}(\vartheta_{2})$ $:=E\left[\log\left(\frac{\phi_{p_{\varphi}}\left(M_{1}^{-1}v_{1};\,(\varphi_{f},0),\,M_{1}^{-1}+\mathrm{diag}(O,\Sigma_{r})\right)}{\phi_{p_{\varphi}}\left(M_{1}^{-1}v_{1};\,(\varphi_{f,0},0),\,M_{1}^{-1}+\Sigma_{0}\right)}\right)\right].$ (3.20)
Then, we assume the following.
@@
  1. (i)
    ${\max_{i\leq N}\left(\lambda_{\min}(M_{i})^{-1}+\lambda_{\min}(\widehat{M}_{i})^{-1}\right)=O_{p}(1)}$.
  2. (ii)
@@
The non-degeneracy condition Assumption 3.5(1) is a technical one. It would be possible to simplify it similarly to the proof of [EguMas19, Lemma 5.2] to handle the inverse of $\widehat{\tau}_{i}$. However, it cannot be done for free, and careful considerations are necessary.
As in [DelGenLar18-2], we introduce the following notation:
$B_{i,T}(\Sigma_{r})$ $=\big(M_{i,T}^{-1}+\mathrm{diag}(0,\Sigma_{r})\big)^{-1},$
$A_{i,T}(\varphi_{f},\Sigma_{r})$ $=B_{i,T}(\Sigma_{r})\big(M_{i,T}^{-1}U_{i,T}-(\varphi_{f},0)\big),$
@@
Finally, let $\mathcal{J}_{2,0}\in\mathbb{R}^{q_{\varphi}}$, where $q_{\varphi}:=p_{\varphi}+p_{\varphi}(p_{\varphi}+1)/2$, denote the covariance matrix associated with the random variable
$\begin{pmatrix}\tau_{1}^{-1}A_{1,T}(\varphi_{f},\Sigma_{r})\\[5.69054pt] \displaystyle{\mathrm{vech}\left(\frac{1}{2}\tau_{1}^{-1}A_{1,T}(\varphi_{f},\Sigma_{r})^{\otimes 2}-B_{1,T}(\Sigma_{r})\right)}\end{pmatrix}$, (3.22)
where $\mathrm{vech}$ denotes the vector-half operator that stacks the lower triangular half into a single vector of length $q_{\varphi}$.
@@
In the current model setup, the random effect parameters in the drift coefficients take the multiplicative form $\tau_{i}\varphi_{i}$, where $\tau_{i}$ and $\varphi_{i}$ are mutually independent. This is not essential. Because of the stepwise nature of our inference procedure, we could flexibly modify this point in several ways; for example, as in [delattre2018parametric] and [DelGenLar18-2], we could incorporate $\varphi_{i}$ instead of $\tau_{i}\varphi_{i}$ with the conditional probability structure $\mathcal{L}(\varphi_{i}|\tau_{i})=N_{p_{\varphi}}(\mu,\tau_{i}\Sigma)$. The modifications are in principle straightforward, and we will not delve into the details here.
@@ 3.3. Connections to existing results
It is worth noting that the convergence rates established in our theoretical analysis are in line with several well-known results in the literature, even though these results were originally obtained under (quasi-)maximum likelihood estimation procedures, which is not exactly the case in our framework. In classical nonlinear mixed-effects models, it is established that, when $N$ and $n$ both tend to infinity fixed effects converge at rate $\sqrt{Nn}$, while parameters governing the distribution of random effects converge at rate $\sqrt{N}$, where $N$ denotes the number of individuals and $n$ the number of observations per individual [nie2007convergence]. In more specialized settings, which correspond to particular cases of our model, it has been shown that, for $T$ fixed and $N,n$ tending to infinity, fixed effects in the diffusion coefficient also converge at rate $\sqrt{Nn}$, while fixed effects in the drift and the parameters of the random effects distributions converge at rate $\sqrt{N}$ [delattre2013maximum, delattre2015estimation, DelGenLar18-2, delattre2018parametric]. These findings are consistent with our results and further supported by the structure of the asymptotic covariance matrix for the estimators of the diffusion parameters, which reveals the same level of precision comparable to that obtained if the individual-level random effects were observed.
@@ 4. Numerical experiments
$\begin{split}Y_{i}(t)&=Y_{i}(0)-\tau_{i}\left(\int_{0}^{t}\frac{\varphi_{f}+\varphi_{r,i}}{\sqrt{1+Y_{i}(s)^{2}}}ds\right)+\sqrt{\tau_{i}}\int_{0}^{t}\exp(\eta\,t/2)dw_{i}(s)\\ \tau_{i}&\sim\textrm{LogNormal}(\alpha,\lambda),\;\varphi_{r,i}\sim_{i.i.d.}N_{1}(0,\sigma_{r}^{2}),\end{split}$
@@
$\begin{split}Y_{i}(t)&=Y_{i}(0)-\tau_{i}\left(\int_{0}^{t}\frac{\varphi_{f}^{(1)}Y_{i}(s)+\varphi_{f}^{(2)}}{\sqrt{1+Y_{i}(s)^{2}}}ds+\int_{0}^{t}\frac{\varphi_{r,i}Y_{i}(s)}{\sqrt{1+Y_{i}(s)^{2}}}\right)ds\,+\\ &\sqrt{\tau_{i}}\int_{0}^{t}\exp(\eta\,\textrm{atan}(Y_{i}(s)))dw_{i}(s)\\ \tau_{i}&\sim\textrm{Weibull}(\alpha,\lambda),\;\varphi_{r,i}\sim_{i.i.d.}N_{1}(0,\sigma_{r}^{2}),\end{split}$
@@
[…]
For each model, data sets are simulated according to different scenarios that combine several numbers of trajectories $N$ (in $\{200,500\}$), several values for $T$ (in $\{5,10\}$) and several time steps $h$ (in $\{0.005,0.001\}$) between consecutive observations leading to different numbers of observations per trajectory $n$. Each observation path is simulated according to its Euler scheme with a very small time step $\delta=0.0001$. In each scenario, $500$ datasets are simulated and the model parameters are estimated using the procedures described above. To illustrate the theoretical results obtained above, the parameters $\theta_{\tau}$ are also estimated from the simulated $\tau_{i}$’s as if they were directly observed. The true parameter values used in this numerical study are as follows.
  • Model 1: $\eta=0.5$, $\alpha=-0.7$, $\lambda=0.7$, $\varphi_{f}=2$, $\sigma_{r}^{2}=1$,
  • Model 2: $\eta=0.5$, $\alpha=1$, $\lambda=0.6$, $\varphi_{f}^{(1)}=2$, $\varphi_{f}^{(2)}=1$, $\sigma_{r}^{2}=1$,
  • Model 3: $\eta=0.5$, $\lambda=1$, $\varphi_{f}^{(1)}=2$, $\varphi_{f}^{(2)}=1$, $\sigma_{r,1}^{2}=1$, $\sigma_{r,1}^{2}=0.5$, $\sigma_{r,1,2}=-0.2$.
The results are depicted in Tables 1-2 (Figure 1) for Model 1, Tables 3-4 (Figure 2) for Model 2 and Tables 5-6 (Figure 3) for Model 3. Each table summarizes the empirical means and standard deviations of the estimates obtained from the $500$ simulated data sets in each of the $(N,n,T)$ configurations.
@@
Across these 3 models, the good results obtained in this simulation study show that the estimation procedure proposed here is sufficiently generic to handle different distributions for random effects in the diffusion coefficient, fixed and random effects in the drift, and a more or less complex covariance structure between random drift effects. The ability to handle any distribution with positive support for the random effect in the diffusion coefficient is a real advance over the state of the art.
Table 1. Simulations under Model 1: empirical mean and empirical standard error (in brackets) of $500$ diffusion parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\eta=0.5$, $\alpha=-0.7$, $\lambda=0.7$. $\widehat{\eta_{y}}$, $\widehat{\alpha_{y}}$ and $\widehat{\sigma_{y}}$ stand for the estimates computed from the trajectories observations, whereas $\widehat{\alpha_{\tau}}$ and $\widehat{\lambda_{\tau}}$ denote the estimates computed from the simulated $\tau_{i}$’s.
@@
Table 2. Simulations under Model 1: empirical mean and empirical standard error (in brackets) of $500$ drift parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\varphi_{f}=2$, $\sigma_{r}^{2}=1$.
$(N,T)$ $n$ $\widehat{\varphi_{f}}$ $\widehat{\sigma_{r}^{2}}$
@@
Table 3. Simulations under Model 2: empirical mean and empirical standard error (in brackets) of $500$ diffusion parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\eta=0.5$, $\alpha=1$, $\lambda=0.6$. $\widehat{\eta_{y}}$, $\widehat{\lambda_{y}}$ and $\widehat{\alpha_{y}}$ stand for the estimates computed from the trajectories observations, whereas $\widehat{\lambda_{\tau}}$ and $\widehat{\alpha_{\tau}}$ denote the estimates computed from the simulated $\tau_{i}$’s.